# Quantifying the noise sensitivity of the Wasserstein metric for images

### v1

@@ A.2 Proof of Theorems
$\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)-T_{\mu_{\epsilon},\nu_{\epsilon}}(x)$ $=\sum_{x\in G_{n}}\mu(x)+{\epsilon}_{\mu}(x)-\nu(x)-{\epsilon}_{\nu}(x)$ (88)
$=0+\sum_{x\in G_{n}}{\epsilon}_{\mu}(x)-{\epsilon}_{\nu}(x),$ (89)
@@
$\sum_{x\in G_{n}}{\epsilon}_{\mu}(x)-{\epsilon}_{\nu}(x)\sim\mathcal{N}(0,2\sigma^{2}N^{2}).$ (92)
@@
$\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)=\sum_{x\in G_{n}}T_{\mu_{\epsilon},\nu_{\epsilon}}(x)\left(1+\frac{\mathrm{O}_{p}\left(\sigma N\right)}{\sum_{x\in G_{n}}T_{\mu_{\epsilon},\nu_{\epsilon}}(x)}\right).$ (93)
@@
$\sum_{x\in G_{n}}T_{\mu_{\epsilon},\nu_{\epsilon}}(x)=\mathrm{O}_{p}(N^{2}).$ (94)
@@
$W_{1}(\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}},\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}})$ $=\sup_{f\in\operatorname{Lip}_{1}}\langle\bar{S}_{\mu_{\epsilon},\nu_{\epsilon}}-\bar{T}_{\mu_{\epsilon},\nu_{\epsilon}},f\rangle$ (95)
$=\sup_{f\in\operatorname{Lip}_{1}}\left\langle\frac{S_{\mu_{\epsilon},\nu_{\epsilon}}}{\sum_{x\in G_{n}}S_{\mu_{\epsilon},\nu_{\epsilon}}(x)}-\frac{T_{\mu_{\epsilon},\nu_{\epsilon}}}{\sum_{x\in G_{n}}T_{\mu_{\epsilon},\nu_{\epsilon}}(x)},f\right\rangle$ (96)

### v2

@@ A.2 Proof of Theorems
$\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)-T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)$ $=\sum_{x\in G_{n}}\mu(x)+\varepsilon_{\mu}(x)-\nu(x)-\varepsilon_{\nu}(x)$ (45)
$=0+\sum_{x\in G_{n}}\varepsilon_{\mu}(x)-\varepsilon_{\nu}(x),$ (46)
@@
$\sum_{x\in G_{n}}\varepsilon_{\mu}(x)-\varepsilon_{\nu}(x)\sim\mathcal{N}(0,2\sigma^{2}N^{2}).$ (49)
@@
$\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)=\sum_{x\in G_{n}}T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)\left(1+\frac{\mathrm{O}_{p}\left(\sigma N\right)}{\sum_{x\in G_{n}}T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}\right).$ (50)
@@
$\sum_{x\in G_{n}}T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)=\mathrm{O}_{p}(N^{2}).$ (51)
@@
$W_{1}(\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}})$ $=\sup_{f\in\operatorname{Lip}_{1}}\langle\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}}-\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}},f\rangle$ (52)
$=\sup_{f\in\operatorname{Lip}_{1}}\left\langle\frac{S_{\mu_{\varepsilon},\nu_{\varepsilon}}}{\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}-\frac{T_{\mu_{\varepsilon},\nu_{\varepsilon}}}{\sum_{x\in G_{n}}T_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)},f\right\rangle$ (53)